Required: 1. Prepare a vertical analysis of Sports Unlimiteds 2018 and 2017 balance sheets. Express each amount as a percent2. Prepare a horizontal analysis of Sports Unlimiteds 2018 balance sheet using 2017 as the base year. (Negative amounts andplease do parts 1 and 2 and show work
Answer #1

Solution:

(1 )Vertical Analysis:

Vertical analysis is also called as Common size analysis. In this method every item of the statement is divided by a common figure, that common figure will be Total of Asset/Lisbilities in case of Balance sheet and Sales in case of Income statement.

For example : Percentage of cash is calculated as

2017= (136500/750000)*100= 18.2%

2018= (232200/900000)*100= 25.8%

Similarly values of all items are calculated. See the attached imageE 18.2 8 19.7 0.9 с D Vertical Analysis of Sports unlimited 2018 2017 3 Items Amounts % Amounts 4 Assets 5 Current Assets: 6

(2) Horizontal Analysis:

Horizontal analysis is also called as Comparative analysis. In this method every item of the statement is compared with the previous year statement, to check the increase or decrease in the value of that item, and the difference is divided by the value of last or base year for calculating percentage .

For example: Cash is compared as follows.

2018 Cash 232200

Less : 2017 Cash 136500

Increase in cash 95700

Percentage increase = (95700/136500)*100= 41.21

Similarly values of all items are calculated. See the attached image below.